answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How do you work out the relative frequency?

The relaive frequency of a particular outcome or event is the number of times the outcome is observed divided by the total number of outcomes observed.

How can probabilities be used in genetics?

Probabilty is to predict the chance of something happening. So probabilties can be used in genetics to predict the chances of things.

For example, if to pea plants mate, one is short and the other is tall, then you could use a Punnett Square to predict what the children would look like. A Punnett Square predicts the possible combination of alleles.

So, in this case (T= tall/dominant and t= short/recessive), if you crossed TT with tt, the offspring would all be tall, or Tt.

If probability of reading newspaper a is .2 probability of reading newspaper b is .16 probability of reading newspaper c is .14.what is the probability that a person selected at random will read none?

The Probability of NOT reading newspaper a is .8 The Probability of NOT reading Newspaper b is .84 The probability of NOT reading Newspaper c is .86 Therefore, .8*.84*.86=0.57792=57.792%

What are numbers that are relatively prime to each other?

Numbers with their Greatest Common Divisor (GCD) as 1 are called to be as relatively prime to each other.

Jacqui is your twin?

Focus: Rainforest Ecosystems, Biomes, Food WebEssential question: How do the animals of the environment of the upperstory, canopy, and forest floor, interact with each other in the food chain based on their presence as a consumer, producer, or decomposer? Rainforests

constant high temperature throughout the year; great quantities of water; humidity Poetry activity

Kinkajou. .- I dreamed I was a soft, fluffy kinkajou hanging from the forest canopy slurping up nectar hungrily.

What is the probability of becoming a soccer player?

well it all depends really infact you ask your self this question first of all you may think your good in your backyard or at school mucking around with younger student and may get a big head and think your good hmm if so advice you to compare your self with people your age if you a star compare them then i would advice to train more and more till your nbetter then people your age also if your not as good as the people your age you can always catch up, i only started football(soccer) 2 years ago and im already playing in super league and indoor soccer well just remember the probability of becoming a soccer player is high because also remmeber (impossible is nothing) think twice before you blow your dream

Suppose you roll 5 number cubes while playing yahtzee. what is the probability of getting a number less than 5 on all of the number cubes?

If there are five dice, there are four numbers less than five, giving each die a 2/3 chance. Over five iterations that is (2/3) to the 5th power or 32/243 or a 13.16 percent chance.

Why is the water bond angle 104.5 less than the CH4 bond angle 109.5?

in water there are two bond pairs and two lone pairs where as in CH4 there are are four bond pairs nad no lone pair.

in ch4 there is only bond pair to bond pair repulsion but in water there are three types of repulsions, lone to lone (greatest repulsion), lone to bond ( lesser repulsion ) and bond to bond ( the least repulsion) , therefore due to the presence of two lone pairs in water the bond pairs are repelled with greater force and they get compressed, reducing the ideal bond angle from 109.5 to 104.5

on the other hand, ch4 has only bond pairs and they dont repel each other that strongly so its angle is greater n its 109.5..

Why were you born a girl?

A girl receives two X sex chromosomes, one from the mother's ovum, and one from the father's sperm cell. Females have the XX sex genotype.

Where did the events in the Bible occur?

The events of the Bible are centered in the Middle East. However, there is mention made of parts of Italy and parts of Greece.

What state quarter comes out next?

The Arizona State Quarter comes out next. New Mexico was the last one and after Arizona, Alaska comes out.

How do you draw spiky hair?

You have to keep practicing, I'm trying to draw spiky hair too (I got floppy hair good). It's hard to explain, but I will explain it as best as I can. You have to make some kind of zig-zag pattern on the head, maybe curve it slightly if you are just drawing the side of the persons head. (LIKE I SAID, ITS HARD TO EXPLAIN) :(

What is the chance of having a baby girl?

It should be 50-50 chance, but recent studies have shown that if you eat vegetables before the conception, there is more chance of it being a girl.

What are the odds of making it in show business?

The exact odds of making it in show business have not actually been calculated. However, those in a position to know what they're talking about say that the odds of getting an occasional role are pretty good, while the odds of being able to make a living from acting are very low. Having another job to pay the bills is vital because there can be long dry spells between acting parts.

'What are the chances your boyfriend is gay when his identical twin brother is gay?

If his identical twin brother is gay, the likelihood that a man is gay it is much higher. Studies have shown that if a man is gay the odds that his brother is gay is 1:2 if they are identical twins; 1:5 if they are fraternal, and 1:20 is they are not twins.

Of course, given that he is your boyfriend, the odds are pretty good that he is straight.

A:

There is an EXTREMELY slim chance identical twins share the same sexuality. Everyone's sexuality is different, and is not (As far as research says) a trait that is passed through genes. Whether someone is gay or not depends on nothing but how that person's brain formed.

IDENTICAL TWINS ARE 99.9% IDENTICAL! They are from the same egg! While there is a slight chance that the other twin is straight, it's not likely. If twins can feel each other's pain both physically or mentally, then I would likely have to say they share the same sexual preference! Good for you, you got two useless hotties! They like men, you're a woman! SCORE! :)

PS: Why don't you just ask him!

I have a standard barrel deluxe checkered Chipmunk .22 rifle 7400 MFG. INC. Medford Ore. What is it value?

The value of a standard barrel deluxe chipmunk .22 rifle 7400 MFG. INC Medford os is definitely worth more than $200. The better condition it is in, the more a consumer will likely pay more.

Are you more fertile after having a baby?

there is no rule on this. it is very individual and varies from woman to woman, day to day and after each pregnancy.

What are the odds of being born?

The odds of being born are astronomicaly low the average man shoots around 20 million sperm. 1 in 20million, but no. Think of it in lower depths the odds of your parents meeting and doing the naughty. Your grandparents being born your great great great great x20 grandparents being born/ meeting eachother. The earth being created. Everything we know. It could be 1 in 9999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999

for all we know.

How thick is the side of a penny?

1.55 mm, which is 0.06102 inches.

The U.S. Mint provides a table of information giving the dimensions, weight, and composition of all current circulation coins. Link below.

What is the chance of getting in the NFL?

The chances are very slim. Even the best players from college can get hosed in the draft, IE Jared Zabransky.

Do you believe there is one or more slot machines in any given casino that are set in the players favor and if so why?

Yes and No. As the saying goes 'the house always wins'. Many slot machines in casinos have a set payout rate; usually in the ranges of about 90-98% (the highest payout rates are in Las Vegas). The 2% that doesn't pay out goes to the casino, they make huge amounts of money from slot machines, even at the highest payout rates.

In terms of them being in players favor many gamblers who have made a living from casinos have always stated that the best chances to make gains are with the slots rather than using a system with roulette or Blackjack. The payouts with the slot machines are on average a lot more than the table games.

When you look at games like Roulette you have a 48% chance of winning on a European Roulette table and 47% chance of winning on the American Table. Compared to a 90-98% payout rate on the slots you can see why a lot of people say the slots are in the players favor.